Determine the molar mass of an unknown nonvolatile solute of nonelectrolyte compound raising the boiling point of 150 g of benzene by 0.500°C.

1
Expert's answer
2022-03-29T01:03:09-0400

Colligative properties:


Boiling point elevation...


∆T = imK


∆T = change in the boiling point = 0.500 degrees


i = van't Hoff factor = 1 for a non electrolyte


m = molal = moles solute/kg solvent = ?


K = boiling point constant for benzene = 2.65º/m (had to look this up since it wasn't provided in the question)


We will solve for m (moles/kg) and since we know there are 0.150 kg (150 g), we an then find moles. We won't be able to find molar mass because we weren't told how many grams of unknown were used. But if you do have that information, you can find molar mass.




Solving for m, we have...


m = ∆T / (i)(K) = 0.5º / (1)(2.65º/m) = 0.189 m = 0.189 moles/kg


0.189 mols/kg x 0.150 kg = 0.0284 moles




If you know the grams used, the molar mass = grams/moles = g/0.0284 = molar mass